### v1

# A Coreset Selection of Coreset Selection Literature: Introduction and Recent Advances

Abstract
Coreset selection targets the challenge of finding a small, representative subset of a large dataset that preserves essential patterns for effective machine learning. Although several surveys have examined data reduction strategies before, most focus narrowly on either classical geometry-based methods or active learning techniques. In contrast, this survey presents a more comprehensive view by unifying three major lines of coreset research, namely, training-free, training-oriented, and label-free approaches, into a single taxonomy. We present subfields often overlooked by existing work, including submodular formulations, bilevel optimization, and recent progress in pseudo-labeling for unlabeled datasets. Additionally, we examine how pruning strategies influence generalization and neural scaling laws, offering new insights that are absent from prior reviews. Finally, we compare these methods under varying computational, robustness, and performance demands and highlight open challenges, such as robustness, outlier filtering, and adapting coreset selection to foundation models, for future research.
Index Terms:
Computer Science, Artificial Intelligence, Coreset Selection, Deep Learning, Survey.
I Introduction
At the heart of coreset selection lies a fundamental research question: How can we identify a small yet representative subset of a massive dataset that preserves the critical structure or distribution needed for accurate learning? This problem was originally proposed by Agarwal et al. [1] to compute the smallest $k$-ball cover for a set of points. In other words, the parameter $k$ is the number of balls that we are allowed to use, and the goal is to minimize how large those balls must be to cover all points. Since then, coreset research has been extended to a wide range of covering problems with profound implications for deep learning [2, 3, 4, 5].
This is particularly relevant in the current era of large-scale datasets and resource-intensive deep learning models [6, 7]. Training on millions of images or text snippets can be prohibitively expensive, both in terms of computation time and memory footprint, as well as energy costs and carbon emissions associated with extensive GPU usage [8, 9, 10]. By focusing on a representative coreset, practitioners can reduce training time, maintain strong performance, and reduce carbon footprints [11, 3]. In real-world scenarios, such as autonomous driving, edge computing, and privacy-constrained applications, strict limits on storage and latency make data-reduction strategies crucial for viable deployment [12, 13, 14].
Despite these early foundational works and comprehensive surveys and benchmarks, the literature now presents a wide range of coreset definitions and data reduction techniques that often differ significantly in formulation and scope [1, 15, 16, 17]. This divergence in definitions and methodologies highlights the evolution of the field and motivates the need for a unifying survey.
In contrast, this survey provides a structured and comprehensive synthesis of coreset selection strategies, categorizing them into three major paradigms: training-free, training-oriented, and blind coreset selection (i.e., approaches that select subsets without leveraging ground-truth labels). We discuss classical selection methods, such as random sampling and geometric approaches; training-informed techniques that leverage loss values, gradient norms, or decision boundaries; and emerging label-free strategies that use clustering or pre-trained models to estimate sample importance. Beyond algorithmic insights, we analyze the theoretical foundations of coreset selection and investigate how pruning strategies affect generalization, neural scaling laws, and dataset redundancy. Empirical comparisons highlight the trade-offs between computational efficiency and model performance across different methods.
Section II lays out definitions of coreset selection and active learning, a closely related field. The connection between these two areas is crucial because many coreset selection strategies naturally extend to active learning by selecting which samples to label. Section III presents training-free selection techniques, such as random sampling and geometry-based methods that still provide a strong baseline for more sophisticated methods. Section IV covers all approaches that fall under the umbrella term of training-oriented methods. They exploit information about training dynamics on the full dataset, such as scoring metrics, decision boundary estimations, submodular functions, gradient matching, and bilevel optimization. Next, Section V complements the previous sections by examining blind coreset selection (label-free data) with pseudo-labeling approaches or vision-language models. To provide practitioners with additional guidelines, Section VI provides more theoretical insights about coreset selection and when to apply which method. Section VII gives an overview of coreset applications to various fields in machine learning. Finally, this work summarizes and points to future directions in Section VIII and Section IX, respectively.
Figure 1: Venn diagram of coreset selection and active learning. Coreset selection typically assumes a labeled dataset, whereas active learning uses small labeled datasets and a large unlabeled pool. Many techniques are shared: both coreset selection and active learning rely on criteria (e.g., uncertainty, diversity, and coverage) to select data subsets.
II Setting and Terminology
In this section, we formalize the coreset selection problem and highlight its close relationship to active learning, as shown in Figure 1. We also briefly touch upon theoretical perspectives that will guide deeper insights in later sections.
II-A Problem Definition: Coreset Selection
We begin with the classic image classification setup, where the training dataset $\mathcal{T}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$ consists of $N$ i.i.d. samples drawn from an underlying data distribution $P$. We define $\mathbf{x}_{i}\in\mathcal{X}$ as the input, and $y_{i}\in\mathcal{Y}$ as its corresponding ground-truth label. Coreset selection aims to derive a subset $\mathcal{S}\subset\mathcal{T}$ with $|\mathcal{S}|\ll|\mathcal{T}|$ such that a model $\theta^{\mathcal{S}}$ trained on $\mathcal{S}$ generalizes as well as a model $\theta^{\mathcal{T}}$ trained on the full dataset $\mathcal{T}$:
$\mathcal{S}^{*}=\mathop{\arg\min}_{\mathcal{S}\subset\mathcal{T}:\frac{|\mathcal{S}|}{|\mathcal{T}|}\leq 1-\alpha}\mathbb{E}_{\mathbf{x},y\sim P}[\mathcal{L}(\mathbf{x},y;\theta^{\mathcal{S}})],$ (1)
where $\alpha\in(0,1)$ is the pruning ratio and $\mathcal{L}$ is a loss function (usually cross-entropy).
Although the above definition is straightforward in principle, finding a suitable $\mathcal{S}$ is challenging [1, 16, 18]. It requires deciding which criteria best measure “importance” or “representativeness” for a given sample $\mathbf{x}$. To put it differently, a quantitative way to enforce an order over the input samples $\mathbf{x}$ is needed to guide the selection process [19, 20, 21]. In the following sections, we explore two high-level paradigms and use the taxonomy shown in Figure 2.
  • Training-Free: Methods that use geometric or statistical properties (e.g., distance in feature space or diversity measures) to select $\mathcal{S}$.
  • Training-Oriented: Methods that exploit partial or full training signals (e.g., loss values, gradients, forgetting events) to identify the most informative samples for the specific model in use.
Figure 2: Taxonomy used in this survey. We differentiate between training-free and training-oriented methods.
II-B Problem Definition: Active Learning
A related research field is active learning, which also aims to reduce the required number of samples from a partially labeled perspective [22, 17, 23]. Instead of selecting a coreset $\mathcal{S}$ from a fully labeled dataset, active learning attempts to determine which unlabeled samples in a large pool $\mathcal{T}_{\mathcal{U}}$ should be sent to an oracle for labeling (e.g., a human annotator or nearest neighbor annotator) [24, 25, 26]. Thus, we aim to reduce the unlabeled dataset $\mathcal{T}_{\mathcal{U}}$, as illustrated in Figure 3.
Formally, consider a large unlabeled dataset $\mathcal{T}_{\mathcal{U}}$ and a small labeled dataset $\mathcal{T}_{\mathcal{L}}$, where $\mathcal{T}_{\mathcal{U}}\cap\mathcal{T}_{\mathcal{L}}=\emptyset$. The goal of active learning is to iteratively augment $\mathcal{T}_{\mathcal{L}}$ by selecting a subset $\mathcal{S}\subset\mathcal{T}_{\mathcal{U}}$ such that $|\mathcal{S}|\ll|\mathcal{T}_{\mathcal{U}}|$ while maximizing the expected model performance on $\mathcal{S}\cup\mathcal{T}_{\mathcal{L}}$. Active learning aims to solve the following optimization problem:
$\mathcal{S}^{*}=\operatorname*{arg\,min}_{\mathcal{S}\subset\mathcal{T}_{\mathcal{U}},\frac{|\mathcal{S}|}{|\mathcal{T}_{\mathcal{U}}|}\leq 1-\alpha}\mathbb{E}_{\mathbf{x},y\sim P}[\mathcal{L}(\mathbf{x},y;\theta^{\mathcal{S}})],$ (2)
where $\theta^{\mathcal{S}}$ is the model trained on $\mathcal{T}_{\mathcal{L}}\cup\mathcal{S}$, $\alpha\in(0,1)$ is the fraction of data left unlabeled, and $\mathcal{L}(\mathbf{x},y;\theta)$ is the model loss (e.g., cross-entropy loss).
In practice, several coreset strategies, especially those focusing on “most informative” or “diverse” examples, can be adapted to active learning settings by simply applying them to unlabeled data and then requesting annotations [27, 28]. Conversely, many established active learning techniques also serve as strong coreset baselines (if the samples are already labeled) [29, 22].
Figure 3: The active learning pipeline. The idea of active learning is to prune a large pool of unlabeled dataset $\mathcal{U}$ and to extend the already existing labeled dataset with essential, newly annotated data.
II-C Coverage
An often overlooked property, especially for earlier coreset selection methods, is coverage [10, 30]. Intuitively, we describe a coreset with good coverage as a set representing the whole range of the dataset distribution, not only the easiest/hardest samples. A well-selected coreset balances the original dataset’s geometric coverage while preserving its structural diversity. Every point in $\mathcal{T}$ is geometrically or in a feature space close to at least one point in a well-selected coreset $\mathcal{S}$. More formally, we can quantify (percentage-wise) coverage $c$ in the following way:
[…]
where $B_{d}(\mathbf{x},r)=\{\mathbf{x}^{\prime}\in X:d(x,x^{\prime})\leq r\}$ is a $r$-radius ball whose center is $\mathbf{x}$, and $d(\cdot,\cdot)$ is a distance metric.
III Training-Free Methods
In this section, we introduce training-free methods, namely, random selection and geometry-based selection.
III-A Random Selection
A straightforward training-free baseline is to select each sample from $\mathcal{T}$ with uniform probability:
$\mathbb{P}\left[(\mathbf{x},y)\in S\right]=\frac{|\mathcal{S}|}{|\mathcal{T}|}=1-\alpha.$ (4)
Despite its simplicity, random sampling remains an important reference in coreset research for two reasons. First, its computational minimalism requires zero overhead beyond the selection step, i.e., no model training, no extra distance computations, etc. Second, in over-parameterized settings or for large, diverse datasets, even a random coreset can perform competitively [31, 15]. Redundant examples in $\mathcal{T}$ indicate that a randomly chosen $\mathcal{S}$ often covers a substantial fraction of the data distribution [32, 33].
Another benefit of random sampling is that it is extremely fast, trivially parallelizable, and unbiased with respect to class or difficulty. However, the latter is also a disadvantage, as it ignores data distribution as well as sample difficulty and tends to become suboptimal when the pruning ratio $\alpha$ increases (i.e., above 90%).
III-B Geometry Based
A fundamental and well-known assumption in machine learning is that samples closer in feature space tend to share similar properties [34, 35]. Geometry-based methods exploit this intuition by identifying redundant or noninformative samples while retaining a subset that effectively represents the original data distribution [36].
Figure 4: Illustration of the iterative selection of Herding. As the angle between the direction to the mean of the real dataset $\mathcal{T}$ and to $\mathbf{x}_{1}$ is smaller than to $\mathbf{x}_{2}$, as well as the distance to $\mathbf{m}_{t-1}$ is higher to $\mathbf{x}_{1}$, $\mathbf{x}_{1}$ will be included for the iteration step $t$, leading to a new coreset mean $\mathbf{m}_{t}$.
III-B1 Herding
Herding is a deterministic coreset selection method that sequentially selects samples that best preserve the statistical properties of the full dataset [37, 38]. Unlike random sampling, Herding greedily selects points to minimize the discrepancy between the coreset mean and the dataset mean in feature space, called moment constraint [39, 40].
Specifically, given a feature mapping $\phi:\mathcal{X}\to\mathbb{R}^{d}$, the moment constraint requires that the empirical mean of the selected coreset approximates that of the full dataset:
$\frac{1}{|\mathcal{T}|}\sum_{\mathbf{x}\in\mathcal{T}}\phi(\mathbf{x})\approx\frac{1}{|\mathcal{S}|}\sum_{\mathbf{x}\in\mathcal{S}}\phi(\mathbf{x}).$ (5)
As shown in Figure 4, Herding iteratively selects the next sample $\mathbf{x}_{t}$ by minimizing the deviation between these two means.
$\mathbf{x}_{t}=\arg\max_{\mathbf{x}\in\mathcal{T}}\langle\phi(\mathbf{x}),\mathbf{m}_{t-1}\rangle,$ (6)
where $\mathbf{m}_{t}=\mathbf{m}_{t-1}+\phi(\mathbf{x}_{t})-\frac{1}{|\mathcal{S}|}\sum_{\mathbf{x}\in\mathcal{S}}\phi(\mathbf{x})$ represents the running mean update and $\langle\mathbf{a},\mathbf{b}\rangle$ is the inner product of two vectors, which measures their alignment.
One benefit of Herding is that it is easy to implement if a suitable embedding $\phi(\cdot)$ is available. While it preserves the global statistics (the mean), its focus may overlook more subtle distribution patterns (like multi-modal clusters). Moreover, it requires iterative updates, which can be costly for very large $\mathcal{T}$.
III-B2 k-Center Greedy
Another popular geometry-based method is k-Center Greedy, which aims to minimize the maximum distance between any point in $\mathcal{T}\setminus\mathcal{S}$ and the “closest” point in $\mathcal{S}$ [22]. Formally, it solves the minimax facility location problem [41]:
$\min_{\mathcal{S}\subset\mathcal{T},\frac{|\mathcal{S}|}{|\mathcal{T}|}=(1-\alpha)}\quad\max_{\mathbf{x}_{i}\in\mathcal{T}\setminus\mathcal{S}}\quad\min_{\mathbf{x}_{j}\in\mathcal{S}}d(\mathbf{x}_{i},\mathbf{x}_{j}),$ (7)
where $d(\mathbf{x}_{i},\mathbf{x}_{j})$ is a predefined distance metric (e.g., Euclidean). The goal is to select $k=(1-\alpha)\cdot{|\mathcal{T}|}$ points that serve as centers, ensuring that every remaining point is close to at least one of these centers. Since solving this problem exactly is NP-hard, an efficient greedy approximation is typically used [22].
The $k$-Center Greedy method can be interpreted as finding a set of representative samples that provide maximal coverage of the dataset under a geometric distance measure. It is commonly used in active learning and coreset selection since it ensures that no part of the dataset is too far from the selected subset [42, 43]. Unlike Herding, which aims to match the dataset embedding mean by enforcing moment constraints, $k$-Center Greedy focuses purely on geometric coverage, targeting a more diverse and well-distributed coreset. However, its strict focus on coverage might not align perfectly with model-specific decision boundaries. Moreover, its computation time is considerably high compared to other methods.
Figure 5: Coreset selection performance of Herding and k-Center Greedy on CIFAR-10 for different feature embeddings (ResNet-18 [44], InceptionNet [45], and VGG [46]).
IV Training-Oriented Methods
As shown in Figure 5, geometry-based methods like Herding and k-Center Greedy heavily rely on the embedding quality of the underlying pre-trained feature extractor. One way to circumvent this problem is to train the model on the target dataset that needs to be pruned.
However, when training on the target dataset is required, one can use more sophisticated methods that also take into account the training dynamics, which fall under the umbrella term of training-oriented methods. By tailoring a subset $\mathcal{S}$ to the actual training dynamics, these methods can often achieve higher accuracy than their training-free counterparts, albeit at the expense of additional computation.
In this section, we present five major families of training-oriented coreset strategies: (i) scoring-based, (ii) decision boundary-based, (iii) submodularity-based, (iv) gradient matching-based, and (v) bilevel optimization-based methods. Throughout, we highlight how each method leverages partial or full training signals, and discuss typical trade-offs regarding efficiency versus precision.
Figure 6: Illustration of scoring-based methods. Usually, they employ a metric derived from multiple training iterations that quantifies the impact of a sample on the training quality (e.g., with forgetting counts, gradient norms, or sensitivity scores). The coreset is then selected by including the highest scores first. Note that the typical power law distribution of scores is only found for sufficiently trained and parameterized neural networks that are used for deriving the scores. Under-parameterized neural networks would lead to a more uniform distribution. Further details can be found in our theoretical remarks in Section VI).
IV-A Scoring-Based
A natural way to incorporate training knowledge is to assign each sample a score that reflects its importance to the learning process. We then construct the coreset by selecting data with the highest scores. Scores can come from forgetting events, gradient norms, or importance sampling criteria, as illustrated in Figure 6.
IV-A1 Forgetting
Capturing forgetting events is beneficial to analyze the stability of sample classifications throughout the training [47]. A forgetting event occurs when a sample, correctly classified at some time step $t$, is later misclassified at $t{\prime}>t$.
Formally, let $y_{i}^{t}=\operatorname*{arg\,max}p(\mathbf{x}_{i};\theta)$ be the predicted label of sample $(\mathbf{x}_{i},y_{i})$ at epoch $t$. A forgetting event is counted whenever:
$\mathbf{1}_{(y_{i}^{t}=y_{i})}=1>\mathbf{1}_{(y_{i}^{t+1}=y_{i})}=0,$ (8)
where $\mathbf{1}$ is the indicator function.
This metric can categorize samples as forgettable (i.e., experiencing one or more forgetting events) or unforgettable (i.e., never forgotten once learned). As a result, we can rank samples based on the number of forgetting events. Empirical studies suggest that unforgettable examples are often well-separated from the decision boundary, which typically indicates redundant information [48, 47, 49]. Forgettable examples, on the other hand, contribute more to refining the decision boundary and generalization. Thus, coreset selection methods based on forgetting take samples with the highest forgetting counts first.
One benefit of utilizing forgetting is that it is easy to implement: just count forgetting events over training epochs. However, it can be highly nondeterministic due to its reliance on factors like optimization schedules (e.g., learning rate decay) and random seeds [48].
IV-A2 GraNd
Another method of scoring samples is to use the Gradient Norm (GraNd), which measures the contribution of each sample to the loss function [50]. Similar to forgetting, we exploit the insight of a training trajectory. Given an epoch $t$, the GraNd score is defined as
$\chi_{t}(\mathbf{x},y)=\mathbb{E}_{\theta_{t}}\|\nabla{\theta_{t}}\mathcal{L}(\mathbf{x},y;\theta_{t})\|_{2},$ (9)
where larger values indicate higher importance, as they indicate a greater force on parameter updates. Thus, we select samples with larger values first.
To reduce the computational cost, an alternative approximation, the Expected L2 Norm (EL2N), replaces the gradient computation with the squared Euclidean norm of the error vector in the output space:
$\chi^{*}_{t}(\mathbf{x},y)=\mathbb{E}_{\theta_{t}}\|p(\mathbf{x};\theta_{t})-y\|_{2}^{2},$ (10)
where $p(\mathbf{x};\theta_{t})$ is the model prediction (i.e., a probability distribution). This approximation avoids backpropagation through the entire network, which was trained beforehand, while still capturing the difficulty of learning a sample.
The benefit of GraNd/EL2N is that it directly relates to the extent to which a sample drives parameter updates. Moreover, they are straightforward ranking metrics once partial training is done. On the other hand, GraNd can be expensive for very large networks or datasets if computed frequently, while EL2N can be imprecise as an approximation in such scenarios. Among all the methods reviewed, we find that GraNd selects the slowest.
IV-A3 Sensitivity
Another importance measure for coreset selection is the sensitivity of a sample, which quantifies its worst-case contribution to the total loss function across model parameters $\theta_{t}$ for multiple epochs $t$:
$s(\mathbf{x},y)=\max_{\theta_{t}}\frac{\mathcal{L}(\mathbf{x},y;\theta_{t})}{\sum_{(\mathbf{x}{\prime},y{\prime})\in\mathcal{T}}\mathcal{L}(\mathbf{x}{\prime},y{\prime};\theta_{t})}.$ (11)
The sampling probability for each sample is then
$\mathbb{P}\left[(\mathbf{x},y)\in S\right]=\frac{s(\mathbf{x},y)}{\sum_{(\mathbf{x}{\prime},y{\prime})\in\mathcal{T}}s(\mathbf{x}{\prime},y{\prime})},$ (12)
i.e., samples with higher sensitivity are more likely to be included in $\mathcal{S}$ as they are critical to preserving structure, while redundant or less informative samples have lower selection probabilities [51, 52].
Importance sampling with sensitivity comes from well-established principles in stochastic optimization [53]. It has a clear interpretation in terms of weighting samples by their expected contribution. However, similar to GraNd, if certain samples have extremely high sensitivity (e.g., outliers), sampling can become heavily skewed toward a small subset, risking overfitting or ignoring coverage of the broader data distribution [54, 55]. Moreover, computing exact sensitivities can be computationally expensive, especially for complex models or high-dimensional data. Also, it can require careful tuning of temperature or smoothing parameters to avoid collapsing on a small subset of high-loss samples [56, 53].
IV-B Decision Boundary Based
Whereas scoring-based approaches track loss signals over time, decision boundary-based methods explicitly estimate how close a sample is to the classification boundary [57, 58]. The assumption underlying these methods is that samples near the boundary tend to be more informative.
Figure 7: Illustration of DeepFool. It calculates for each data sample a perturbation $\mathbf{r}$ and selects the samples with the shortest perturbations first (minimal norm). Thus, it prioritizes data near the decision boundary.
IV-B1 DeepFool
DeepFool is a geometry-based method that approximates the distance of a sample to the decision boundary by leveraging adversarial perturbations [27, 59]. The key idea is that samples near the decision boundary carry more information for classification than those farther away, making them ideal for coreset selection, as illustrated in Figure 7.
Given a trained classifier $p(\mathbf{x};\theta)$, DeepFool iteratively finds the smallest perturbation $\mathbf{r}$ needed to move a sample $\mathbf{x}$ across the decision boundary:
$\mathbf{x}{\prime}=\mathbf{x}+\mathbf{r},\text{ where }\mathbf{r}=\operatorname*{arg\,min}_{\mathbf{r}{\prime}}\|\mathbf{r}{\prime}\|_{2}$
[…]
For coreset selection, we prioritize samples with the smallest perturbation magnitude $\|\mathbf{r}\|_{2}$, as they are closest to the decision boundary and, thus, more informative. The DeepFool-based method ranks all samples by $\|\mathbf{r}\|_{2}$ and selects the top k with the smallest perturbation norms.
In contrast to training-free geometry-based methods that do not account for the actual learned boundary, DeepFool often yields improved performance. However, it requires repeated adversarial steps, which can be expensive and sensitive to the hyperparameters of adversarial routines (e.g., step size).
IV-B2 Contrastive Active Learning
Contrastive Active Learning (CAL) proposes an acquisition strategy that balances uncertainty sampling and coverage sampling by selecting contrastive examples, i.e., data points that are similar in feature space but yield highly divergent model predictions [60, 61, 62]. This approach ensures that selected examples contribute maximally to refining the decision boundary of the model.
Specifically, CAL defines a contrastive pair $(\mathbf{x}_{i},\mathbf{x}_{j})$ based on two criteria: Feature similarity and predictive discrepancy. Regarding the first, CAL is optimized to map the data points to be close in the feature space:
$d(\Phi(\mathbf{x}_{i}),\Phi(\mathbf{x}_{j}))<\epsilon$ (14)
where $\Phi(\cdot)$ is an encoder mapping input samples to a shared feature space, $d(\cdot,\cdot)$ is a distance metric, and $\epsilon>0$ is a small threshold. With respect to predictive discrepancy, the model should assign significantly different predictive probabilities to the two examples:
$D_{\text{KL}}(p(y|\mathbf{x}_{i})\,||\,p(y|\mathbf{x}_{j}))\to\infty,$ (15)
where $D_{\text{KL}}$ denotes the Kullback-Leibler (KL) divergence. Intuitively, contrastive examples are samples that lie near the decision boundary, where the model is uncertain. CAL prioritizes these samples for the coreset.
One benefit of CAL is that it balances data diversity (i.e., closeness in feature space) with uncertainty (i.e., mismatched predictions). Nevertheless, it also requires computing pairwise distances and KL divergences, which can be costly at scale. Similar to geometry-based, training-free approaches, the performance of CAL also depends heavily on the quality of the learned embedding.
Figure 8: Example of a submodular function where $A=\{\mathbf{x}_{1}\}\subset B=\{\mathbf{x}_{1},\mathbf{x}_{2}\}$: $B_{d}(\mathbf{x},r)=|\{\mathbf{x}^{\prime}\in X:d(x,x^{\prime})\leq r\}|$, where $B_{d}(\mathbf{x},r)$ is a $r$-radius ball whose center is $\mathbf{x}$. Naturally, a smaller set of centers has a higher chance of encapsulating more distinguished points.
IV-C Submodularity Based
Submodularity is a fundamental property of set functions that captures the principle of diminishing returns, which makes it especially interesting for coreset selection [63, 64, 65, 29, 66].
A function $f:2^{V}\to\mathbb{R}$ defined over a ground set $V$ is submodular if, for any subsets $A\subseteq B\subseteq V$ and any element $j\in V\setminus B$ , the following holds:
$f(A\cup\{j\})-f(A)\geq f(B\cup\{j\})-f(B).$ (16)
This property, known as the diminishing returns condition, means that adding an element to a smaller set provides a greater marginal gain than adding it to a larger set. Figure 8 shows an example. Intuitively, submodular functions model concepts like diversity, coverage, and representativeness.
In general, our goal in coreset selection is to maximize a submodular function:
$\mathcal{S}^{*}=\operatorname*{arg\,max}_{\mathcal{S}\subseteq V,\frac{|\mathcal{S}|}{|\mathcal{T}_{u}|}\leq 1-\alpha}f(\mathcal{S}).$ (17)
Common submodular functions in coreset selection are Facility Location (maximizing coverage) with
$f(\mathcal{S})=\sum_{\mathbf{x}\in\mathcal{T}}\max_{\mathbf{x}{\prime}\in\mathcal{S}}\text{sim}(\mathbf{x},\mathbf{x}{\prime}),$ (18)
where $\text{sim}(\mathbf{x},\mathbf{x}{\prime})$ is a similarity metric (e.g., cosine similarity) and Graph Cut (minimizing redundancy) with
$f(\mathcal{S})=\sum_{\mathbf{x}\in\mathcal{S}}\sum_{\mathbf{x}{\prime}\in\mathcal{S}}w(\mathbf{x},\mathbf{x}{\prime}),$ (19)
where $w(\mathbf{x},\mathbf{x}{\prime})$ represents a similarity weight.
However, this problem is NP-hard, meaning that finding the exact optimal subset $\mathcal{S}^{*}$ is computationally infeasible for large datasets. Luckily, submodular functions exhibit a key optimization advantage: they can be efficiently maximized using a greedy algorithm, which provides a $(1-1/e)\approx 63\%$approximation guarantee for cardinality-constrained maximization [67]. This guarantee means that if $f(\mathcal{S}^{*})$ is the best possible value, then the greedy-selected subset $\mathcal{S}_{\text{greedy}}$ satisfies:
$f(\mathcal{S}_{\text{greedy}})\geq(1-1/e)\cdot f(\mathcal{S}^{*}).$ (20)
This tells us a few things: 1) Greedy methods give strong approximations without needing brute-force optimization. 2) The greedy algorithm will never be worse than about 63% of the best possible solution. 3) Lazy greedy techniques can be employed to speed up computation significantly [68].
A natural arising question is: Why is it not 100% optimal? The reason we cannot achieve exactly $f(\mathcal{S}^{*})$ with a greedy method is that each step picks the locally best option without accounting for global interactions in the dataset. This is known as greedy suboptimality. However, for most practical submodular maximization problems (including coreset selection), the greedy algorithm with a $(1-1/e)$ guarantee is a well-established strong baseline.
IV-C1 FASS
Filtered Active Submodular Selection (FASS) is a method designed to improve active learning by combining uncertainty sampling with submodular optimization [29]. The key idea is to filter uncertain examples first and then select the most representative subset using submodular optimization.
Consider a dataset $\mathcal{T}$ consisting of labeled ($\mathcal{T}_{l}$) and unlabeled ($\mathcal{T}_{u}$) samples: $\mathcal{T}=\mathcal{T}_{l}\cup\mathcal{T}_{u}$. FASS first selects a candidate pool $\mathcal{C}$ of uncertain examples from $\mathcal{T}_{u}$. A typical uncertainty measure is entropy:
$H(\mathbf{x})=-\sum_{c}p(y_{c}|\mathbf{x};\theta)\log p(y_{c}|\mathbf{x};\theta),$ (21)
where $p(y_{c}|\mathbf{x};\theta)$ is the model prediction probability of class c. The candidate pool $\mathcal{C}$ is then created from the top $m$ most uncertain samples:
$\mathcal{C}=\operatorname*{arg\,max}_{\mathcal{C}\subseteq\mathcal{T}_{u},|\mathcal{C}|=m}\sum_{\mathbf{x}\in\mathcal{C}}H(\mathbf{x})$ (22)
Given the candidate set $\mathcal{C}$, we select a batch $\mathcal{S}$ of size $(1-\alpha)\cdot|\mathcal{T}|$ by maximizing a submodular function $f$:
$\mathcal{S}^{*}=\operatorname*{arg\,max}_{\mathcal{S}\subseteq\mathcal{C},\frac{|\mathcal{S}|}{|\mathcal{T}|}=(1-\alpha)}f(\mathcal{S}).$ (23)
The key benefit of FASS is that it reduces the computational cost by restricting submodular selection to uncertain points, potentially merging the best of both worlds - uncertainty and coverage. However, the computational costs are still considerably higher than those of other coreset selection methods.
IV-C2 PRISM
In a similar vein, PRISM (Parameterized Submodular Information Measures) builds on submodular information measures, namely CG, MI, and CMI [69]. In the following, we will explain each measure in detail. The submodular Conditional Gain (CG) measures the gain in function value when adding a set $\mathcal{S}$ to a validation set $\mathcal{V}=\mathcal{T}\setminus\mathcal{S}$:
$f(\mathcal{S}|\mathcal{V})=f(\mathcal{S}\cup\mathcal{V})-f(\mathcal{V}).$ (24)
CG is used to quantify how different a selected set $\mathcal{S}$ is from $\mathcal{V}$. On the other hand, the Submodular Mutual Information (MI) captures the similarity between sets $\mathcal{S}$ and $\mathcal{T}$:
$I_{f}(\mathcal{S};\mathcal{T})=f(\mathcal{S})+f(\mathcal{T})-f(\mathcal{S}\cup\mathcal{T})$ (25)
MI encourages selecting elements similar to $\mathcal{T}$. By combining both, we obtain Submodular Conditional Mutual Information (CMI), which jointly models both the similarity to $\mathcal{T}$ and dissimilarity from $\mathcal{V}$:
$I_{f}(\mathcal{S};\mathcal{T}|\mathcal{V})=f(\mathcal{S}|\mathcal{V})+f(\mathcal{T}\cup\mathcal{V})-f(\mathcal{S}\cup\mathcal{T}\cup\mathcal{V})$ (26)
Thus, the overall optimization goal for a subset is
$\mathcal{S}^{*}=\operatorname*{arg\,max}_{\mathcal{S}\subseteq\mathcal{T},\frac{|\mathcal{S}|}{|\mathcal{T}_{u}|}\leq 1-\alpha}I_{f}(\mathcal{S};\mathcal{T}|\mathcal{V}).$ (27)
PRISM shares the advantages and disadvantages of FASS. Still, the CMI-based framework unifies the notion of coverage (similar to facility location) with a penalty for being too similar to existing sets.
IV-C3 SIMILAR
SIMILAR (Submodular Information Measures Based Active Learning In Realistic Scenarios) extends submodular coreset selection by leveraging Submodular Mutual Information (SMI) and Submodular Conditional Mutual Information (SCMI) to select diverse and informative subsets [28]. The framework unifies uncertainty sampling, diversity-based selection, and coverage-based selection within a single mathematical formulation.
SMI quantifies the mutual dependence between a selected coreset $\mathcal{S}$ and a reference set $\mathcal{Q}$ (which could be the full dataset, rare-class samples, or out-of-distribution data) using:
$I_{f}(\mathcal{S};\mathcal{Q})=f(\mathcal{S})+f(\mathcal{Q})-f(\mathcal{S}\cup\mathcal{Q}),$ (28)
where $f(\cdot)$ is a submodular function modeling representativeness or diversity.
To further refine the selection process, SIMILAR introduces SCMI, which penalizes redundancy by conditioning on an already selected set $\mathcal{V}$:
$I_{f}(\mathcal{S};\mathcal{Q}|\mathcal{V})=f(\mathcal{S}|\mathcal{V})+f(\mathcal{Q}|\mathcal{V})-f(\mathcal{S}\cup\mathcal{Q}|\mathcal{V}).$ (29)
This enables a more context-aware selection strategy, ensuring that selected samples are informative relative to both the dataset $\mathcal{Q}$ and the existing knowledge in $\mathcal{V}$.
The coreset selection follows:
$\mathcal{S}^{*}=\operatorname*{arg\,max}_{\mathcal{S}\subseteq\mathcal{T},\frac{|\mathcal{S}|}{|\mathcal{T}|}\leq 1-\alpha}I_{f}(\mathcal{S};\mathcal{Q}|\mathcal{V}),$ (30)
which allows for adaptive batch selection, ensuring diversity and non-redundancy.
SIMILAR provides a generalized coreset selection framework applicable to rare-class selection, out-of-distribution data filtering, and uncertainty-aware sampling. By leveraging lazy greedy optimization, it maintains a $(1-1/e)$ approximation guarantee while being more scalable than traditional submodular methods. However, its reliance on computing large similarity matrices increases its complexity, making it less efficient for very large datasets unless approximate submodular maximization techniques are used.
IV-D Gradient Matching Based
In contrast to previous approaches, gradient matching focuses on reconstructing the total gradient of the full dataset using only a subset of samples. This ties subset selection directly to the training optimization path.
IV-D1 CRAIG
CRAIG (Coresets for Accelerating Incremental Gradient Descent) is a coreset selection method designed to select a subset of data points whose weighted gradients closely approximate the gradient of the full dataset [70]. The optimization target is to find the smallest subset $\mathcal{S}$ and associated step sizes $\gamma_{j}>0$ such that:
$\max_{\theta\in W}\left\|\sum_{i\in\mathcal{T}}\nabla_{\theta}\mathcal{L}(\mathbf{x}_{i},y_{i};\theta)-\sum_{j\in\mathcal{S}}\gamma_{j}\nabla_{\theta}\mathcal{L}(\mathbf{x}_{j},y_{j};\theta)\right\|\leq\epsilon.$ (31)
Directly solving this optimization is computationally infeasible. However, the problem can be transformed into a submodular facility location function, allowing a greedy algorithm to efficiently approximate the optimal subset. We can incrementally build $\mathcal{S}$ by iteratively adding the sample that maximizes the marginal gain in gradient approximation. Each sample $\mathbf{x}_{j}\in\mathcal{S}$ is assigned a weight $\gamma_{j}$, which is determined based on the number of samples in $\mathcal{T}$ that are closest to it in the gradient space.
The selected subset matches the full-dataset gradient well, accelerating training while maintaining performance. Yet, it requires partial training (or iteratively updated gradients) to identify $\gamma_{j}$ weights, which is an additional overhead.
IV-D2 GradMatch
Similar in spirit is GradMatch [71]. However, instead of ranking samples in isolation, GradMatch iteratively constructs a coreset by minimizing the gradient reconstruction error, ensuring that the subset provides a globally representative gradient direction. Mathematically, the objective is to minimize the gradient matching error:
$\mathcal{E}(\mathcal{S})=\lambda\left\|\gamma\right\|^{2}+$ (32)
$\left\|\sum_{\mathbf{x}_{i}\in\mathcal{T}}\nabla_{\theta}\mathcal{L}(\mathbf{x}_{i},y_{i};\theta)-\sum_{\mathbf{x}_{j}\in\mathcal{S}}\gamma_{j}\nabla_{\theta}\mathcal{L}(\mathbf{x}_{j},y_{j};\theta)\right\|^{2},$
GradMatch employs Orthogonal Matching Pursuit (OMP) [72] to iteratively select the subset $\mathcal{S}$ by maximizing the gradient similarity. At each step $t$, the algorithm selects the sample $\mathbf{x}_{t}$ that best reduces the gradient matching error:
$\mathbf{x}_{t}=\operatorname*{arg\,max}_{\mathbf{x}_{i}\in\mathcal{T}\setminus\mathcal{S}}\left\|\langle\nabla_{\theta}\mathcal{L}(\mathbf{x}_{i},y_{i};\theta),\mathbf{r}_{t-1}\rangle\right\|,$ (33)
where $\mathbf{r}_{t}$ is the residual gradient:
$\mathbf{r}_{0}$ $=\sum_{\mathbf{x}_{i}\in\mathcal{T}}\nabla_{\theta}\mathcal{L}(\mathbf{x}_{i},y_{i};\theta),\text{ and}$
$\mathbf{r}_{t}$ $=\mathbf{r}_{t-1}-\gamma_{t}\nabla_{\theta}\mathcal{L}(\mathbf{x}_{t},y_{t};\theta).$ (34)
GradMatch achieves direct alignment with optimizer updates and often yields impressive compression without accuracy loss. Nevertheless, it comes with potentially high costs for large datasets, as each step involves searching $\mathcal{T}\setminus\mathcal{S}$. In practice, GradMatch shows sensitivity to hyperparameters (e.g., number of OMP iterations) [71, 72].
IV-E Bilevel Optimization Based
Bilevel optimization frameworks explicitly treat coreset selection as an “outer” problem that optimizes model performance on a validation or labeled set, while the “inner” problem performs standard training. Although more computationally complex, these approaches can produce subsets tuned for generalization rather than simply reconstructing training gradients [73, 74, 75].
IV-E1 RETRIEVE
The RETRIEVE method [76] extends gradient-based coreset selection to the Semi-Supervised Learning (SSL) setting [77], where both labeled and unlabeled data must be leveraged effectively. Unlike GradMatch, which formulates coreset selection as a gradient matching problem in a fully supervised setting, RETRIEVE introduces a bilevel optimization framework that selects the most informative unlabeled samples to improve generalization when labeled data are scarce.
Let $\mathcal{T}=\mathcal{T}_{l}\cup\mathcal{T}_{u}$ denote a dataset consisting of a labeled set $\mathcal{T}_{l}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N_{l}}$ and an unlabeled set $\mathcal{T}_{u}=\{\mathbf{x}_{i}\}_{i=1}^{N_{u}}$. The goal of coreset selection in SSL is to construct a compact subset $\mathcal{S}\subset\mathcal{T}_{u}$ such that training on $\mathcal{T}_{l}\cup\mathcal{S}$ results in a model with a similar or superior generalization to training on the full dataset.
Formally, RETRIEVE optimizes a bilevel objective:
$\mathcal{S}^{*}=\arg\min_{\mathcal{S}\subset\mathcal{T}_{u},\frac{|\mathcal{S}|}{|\mathcal{T}_{u}|}\leq 1-\alpha}\sum_{(\mathbf{x},y)\in\mathcal{T}_{l}}\mathcal{L}(\mathbf{x},y;\theta^{*}(\mathcal{S})),$ (35)
where $\mathcal{L}(\mathbf{x},y;\theta)$ is the supervised loss (e.g., cross-entropy), and the inner optimization defines the optimal model parameters $\theta^{*}$ as:
$\theta^{*}(\mathcal{S})=\arg\min_{\theta}\left(\sum_{(\mathbf{x},y)\in\mathcal{T}_{l}}\mathcal{L}(\mathbf{x},y;\theta)+\lambda\sum_{\mathbf{x}\in\mathcal{S}}\mathcal{L}_{u}(\mathbf{x};\theta)\right).$ (36)
Here, $\mathcal{L}_{u}(\mathbf{x};\theta)$ represents the unsupervised loss applied to unlabeled samples, typically a consistency regularization term such as Mean Squared Error (MSE) between different augmentations.
Solving the bilevel optimization directly is computationally intractable. Instead, RETRIEVE employs a one-step gradient-based approximation to estimate the impact of selecting an unlabeled sample $\mathbf{x}\in\mathcal{T}_{u}$ on the labeled set’s loss:
$s(\mathbf{x})=\left\|\nabla_{\theta}\sum_{(\mathbf{x},y)\in\mathcal{T}_{l}}\mathcal{L}(\mathbf{x},y;\theta)\right\|_{2}-\lambda\left\|\nabla_{\theta}\mathcal{L}_{u}(\mathbf{x};\theta)\right\|_{2}.$ (37)
The coreset $\mathcal{S}$ is then selected by ranking samples according to $s(\mathbf{x})$, ensuring that unlabeled samples with high influence on the labeled loss are prioritized.
RETRIEVE is explicitly designed for SSL, leveraging partial labels, and may be too impractical in fully labeled scenarios.
IV-E2 GLISTER
Another intriguing bilevel optimization method is GLISTER (GeneraLIzation-based data Subset selecTion for Efficient and Robust learning), which selects a subset that maximizes log-likelihood performance on a held-out validation set rather than just the training set [78]. This shift prioritizes generalization while maintaining efficiency.
Mathematically, GLISTER optimizes as follows:
$\operatorname*{arg\,max}_{\mathcal{S}\subseteq\mathcal{T}\setminus\mathcal{V},\frac{|\mathcal{S}|}{|\mathcal{T}\setminus\mathcal{V}|}\leq 1-\alpha}\mathcal{L}_{\mathcal{V}}\left(\mathcal{V},\operatorname*{arg\,max}_{\theta}\mathcal{L}_{\mathcal{T}\setminus\mathcal{V}}(\mathcal{S},\theta)\right),$ (38)
where $\mathcal{V}\subset\mathcal{T}$ is the validation set, $\mathcal{L}_{\mathcal{V}}$ the validation loss, and $\mathcal{L}_{\mathcal{T}\setminus\mathcal{V}}$ is the training loss on the remaining data that are candidates for the coreset.
The key advantage of GLISTER is its focus on “held-out” performance that can improve real-world robustness (e.g., fewer overfitting concerns). However, it requires setting aside a validation set $\mathcal{V}$, reducing the total data for training and, thereby, potentially removing beneficial samples for the final coreset.
Figure 9: Comparison of method runtimes, selection plus training, in minutes (tested on RTXA6000), sorted in ascending order. The histogram highlights the varying computational costs of the different selection strategies, with GraNd being the most time-intensive and random being the fastest.
IV-F Computation Time
Figure 9 illustrates the computation time for various coreset methods, sorted in ascending order. This comparison shows the trade-offs between computational efficiency and method complexity, which must be carefully considered when choosing a suitable approach for large-scale applications. Notably, Random, Forgetting, GradMatch, Herding, and Craig exhibit the lowest computation times, with k-Center Greedy and DeepFool requiring significantly more time. The most computationally expensive method, GraNd, stands out with a runtime of 125 minutes (for ImageNet-1K, it takes 1.5 days on a single RTXA6000 GPU). Methods like CAL, Glister, GraphCut, and Facility Location fall in the mid-range, balancing computational cost and selection quality.
TABLE I: Overview of coreset selection methods discussed in this review.
Category Subcategory Method Comments
Training-free Random Random Selection Simple and fast. Poor performance at high pruning rates.
Geometry-based K-Center Greedy Finds maximal coverage under a geometric measure. Suitable for clustering tasks. Computationally expensive for larger datasets.
Geometry-based Herding Greedily selects samples to minimise distance between coreset and dataset centroids. Fast and easy to implement. Poor performance on complex data distributions.
Training-based Score-based Forgetting Counts forgetting events during training and picks samples with highest counts. Easy to implement. Highly non-deterministic.
GraNd Picks samples with maximal contribution to the loss function during training. Easy implementation. Expensive and slow for large datasets.
Decision Boundary-based DeepFool [27, 59] Iteratively finds the smallest perturbation to move samples across a decision boundary. Picks samples with smallest values. Expensive owing to repeated adversarial steps.
CAL Balances uncertainty and diversity by picking contrastive samples. Computationally intensive; Dependent on embedding quality.
Submodularity-based FASS [29] Combines informativeness and representativeness. Filters out uncer Computationally expensive.
PRISM [69] Adaptable to different information measures; robust to noise. Requires careful parameter tuning. May not be efficient for very large datasets.
SIMILAR [28] Captures both informativeness and diversity. Computationally expensive due to submodular optimization. Requires labeled data.
Gradient Matching-based CRAIG [70]. Efficiently accelerates gradient descent. Reduces computational cost. Limited to specific optimization algorithms. May not generalize well to all tasks.
Gradmatch [71] Matches gradients of the coreset to the full dataset. Effective for neural networks. Requires computing gradients for the entire dataset. Sensitive to hyperparameters.
Bilevel Optimization-based RETRIEVE [76] Iteratively refines coreset based on model predictions. Computationally intensive due to iterative refinement. Requires model retraining.
GLISTER [78] Generalizes well across different models; Robust to noise and outliers. Requires careful tuning of generalization metrics. May not be efficient for very large datasets.
V Blind Coreset Selection
Traditional coreset selection methods rely on labeled data to estimate sample importance via loss values, gradient norms, or decision boundary distances. In contrast, blind coreset selection methods aim to approximate these training dynamics without ground-truth annotations.
Figure 10: Illustration of ELFS. First, it embeds the dataset in a feature space embedding and applies deep clustering to assign pseudo-labels. Next, it applies coreset selection based on scoring (forgetting, gradient norms, and sensitivity).
V-A ELFS
The Estimation of Label-Free Sample Scores (ELFS) framework [79] achieves label-free coreset selection by leveraging pseudo-labeling through deep clustering [80] to approximate informative selection criteria. More formally, given an unlabeled dataset $\mathcal{T}=\{\mathbf{x}_{i}\}_{i=1}^{N}$, ELFS estimates pseudo-labels $\hat{y}_{i}$ by clustering feature embeddings extracted from a pre-trained network $\phi$:
$\mathbf{z}_{i}=\phi(\mathbf{x}_{i})\quad\forall\mathbf{x}_{i}\in\mathcal{T}.$ (39)
The pseudo-label assignment is then performed via deep clustering:
$\hat{y}_{i}=\arg\max_{k}p(y=k|\mathbf{z}_{i}),$ (40)
where $p(y|\mathbf{z}_{i})$ is estimated using a clustering algorithm such as teacher ensemble-weighted mutual information [81]. Once pseudo-labels are assigned, ELFS estimates training dynamics metrics, such as forgetting events, gradient norms, and sample sensitivity.
The whole procedure is visualized in Figure 10. By leveraging pseudo-labels, it mimics training-informed, well-known coreset selection methods (e.g., forgetting events, gradient norms) without explicit supervision. However, the effectiveness of ELFS is highly dependent on how well the clustering algorithm assigns pseudo-labels. This problem becomes pronounced in high-dimensional embeddings, where feature separability is challenging, leading to poor clustering performance.
Figure 11: Coreset selection performances on CIFAR-10. Methods grouped by our taxonomy a colored the same, with dashed lines indicating the different methods within the same category. A general observation is that methods that are good for high pruning ratios ($\alpha$) fail to maintain the performance at low pruning ratios and vice versa. We can see a transition between top-performing methods between 5% and 30% of the remaining dataset size. Concrete values can be found in Table II.
V-B ZCore
Zero-shot Core (ZCore) circumvents the need for model-specific supervision or clustering by employing pre-trained Vision-Language Models (VLMs) to encode feature representations [82]. Given an unlabeled dataset $\mathcal{T}$, the method first embeds each sample into a feature space using a pre-trained foundation model $\phi$ (similar to geometry-based methods):
$\mathbf{z}_{i}=\phi(\mathbf{x}_{i})\quad\forall\mathbf{x}_{i}\in\mathcal{T}$ (41)
Next, ZCore applies a coverage-based selection strategy inspired by k-Center Greedy and submodular maximization. The goal is to identify a subset such that every point in $\mathcal{T}$ is close to at least one representative sample in $\mathcal{S}$, which is expressed as a score function $s$:
$\mathcal{S}=\operatorname*{arg\,max}_{\mathcal{S}\subset\mathcal{T},\frac{|\mathcal{S}|}{|\mathcal{T}|}\leq 1-\alpha}\sum_{\mathbf{x}_{i}\in\mathcal{S}}s(\mathbf{x}_{i}).$ (42)
The key advantage of ZCore is that by eliminating the dependence on model-specific training, it can be applied to any unlabeled dataset without additional computational overhead. However, its reliance on pre-trained representations introduces potential domain bias (similar to other training-free methods). Additionally, while ZCore ensures geometric coverage, it does not explicitly account for task-specific importance, which may limit performance in highly specialized applications.
VI Theoretical Remarks
A fundamental question in coreset selection, and through the lens of the broader field of dataset pruning, is whether a smaller, carefully chosen subset $\mathcal{S}$ can match or even surpass the performance of the full dataset $\mathcal{T}$. This challenge ties into neural scaling laws [83, 84, 85, 86], which describe how the generalization error $\epsilon(\mathcal{T})$ typically follows a power-law decay:
$\epsilon(\mathcal{T})\propto|\mathcal{T}|^{-\beta}$ (43)
where $\beta$ is a dataset-dependent exponent. This suggests that collecting more data improves model performance but with diminishing returns. However, coreset selection and data pruning in general disrupt this traditional scaling, as they introduce selection bias by favoring the most informative samples [87, 88]. Instead of following the standard power law, a well-pruned dataset can achieve:
$\epsilon(\mathcal{S})\propto|\mathcal{S}|^{-\beta{\prime}},$ (44)
where $\beta{\prime}>\beta$.
This means that with the right subset selection, fewer samples can achieve better generalization than an equally large randomly selected subset. The effective information content of a dataset can be defined as:
$I_{\text{eff}}(\mathcal{T})=\sum_{(\mathbf{x},y)\in\mathcal{T}}w(\mathbf{x},y)H(\mathbf{x},y),$ (45)
where $w(\mathbf{x},y)$ is a weighting function that adjusts for redundancy, and $H(\mathbf{x},y)$ is the sample entropy [89]. By filtering out samples with the lowest information contribution, ensuring that only the most representative data remains.
Sorscher et al. [90] provide a deeper analysis of how pruning affects generalization. Their study reveals that the optimal subset selection strategy depends on the size of the original dataset:
  • For small datasets ($|\mathcal{T}|$ is small), keeping easy (high-margin) samples is better. These samples provide coarse-grained information, which prevents overfitting and ensures stable learning.
  • For large datasets ($|\mathcal{T}|$ is large), keeping hard (low-margin) samples is more effective. These samples refine the decision boundary and contribute to fine-grained generalization.
Moreover, their findings indicate that the pruning ratio should increase as the dataset size increases. This is because, in large datasets, much of the information is redundant, and a smaller fraction of harder examples is sufficient for optimal learning.
Figure 12: Coreset selection performances on CIFAR-10 with 10% corrupted labels, which means that 10% of samples have been reassigned to new class labels [91]. Methods grouped by our taxonomy a colored the same, with dashed lines indicating the different methods within the same category. A general observation is that most methods experience a sharp drop in mid-range pruning ratios (10%-50%), especially geometry-based, gradient matching-based methods, and GraNd (gradient norm). Forgetting, CAL, and GraphCut remain more stable in comparison to the other methods. Concrete values can be found in Table III.
For coreset selection, we see indications of these insights by looking at the performances of different methods on CIFAR-10, as shown in Figure 11 and with concrete values in Table II. Methods emphasizing easy samples are outperforming for higher pruning ratios (smaller dataset size) and underperforming for lower ratios and vice versa (see take-over points at around 10% dataset size). However, Zheng et al. [10], and others [92], partially disagree as they found out that always keeping a small fraction of high-density (redundant) is further boosting the performance alongside a hardest-first sampling strategy.
A crucial takeaway from their work is that pruning can break traditional power law scaling. While random subsampling follows the expected decay $\beta$, a well-pruned dataset can achieve an improved scaling exponent $\beta{\prime}$, leading to exponential improvements in generalization error. This means that dataset pruning is not just about compression - it is a powerful tool for enhancing model performance by maximizing information efficiency.
VI-A Outliers
Another interesting aspect of coreset selection, often overlooked, is the role of outliers [93, 94]. The prior handling of outliers can significantly improve the performance of any coreset selection method [95]. Outliers can either provide valuable decision boundary information or introduce noise and perturbations during the training.
We provide in Figure 12 and Table III additional robustness experiments to highlight the importance of outliers. In this experiment, we corrupted 10% of the labels, i.e., reassigning 10% of samples to new class labels [91]. The selection of corrupted labels was repeated during each coreset sampling procedure.
As a result, one can see that most sampling methods get negatively affected (red), especially geometry-based, gradient-matching-based methods, and GraNd (gradient norm). Interestingly, forgetting, CAL, and GraphCut remain more stable in comparison to the other methods. Furthermore, CAL confirms Equation 44, as its $\alpha=10\%$ pruned dataset outperforms the training on the full (corrupted) dataset. Also, forgetting and GraphCut were able to achieve better results for some cases (green).
As this experiment highlights, future work could further explore automated ways prior to coreset selection to detect and filter outliers dynamically. For instance, Zheng et al. [10] leveraged the area under margin [96] for their coverage-based coreset selection, which captures how far a sample is from the decision boundary over multiple training iterations. Other innovations might include classical outlier detection applications.
TABLE II: Coreset selection performances on CIFAR10 with randomly initialized ResNet-18 [44] models. Values derived from DeepCore [15].
Fraction $(1-\alpha)$ 0.1% 0.5% 1% 5% 10% 20% 30% 40% 50% 60% 90% 100%
Random 21.0 $\pm$ 0.3 30.8 $\pm$ 0.6 36.7 $\pm$ 1.7 64.5 $\pm$ 1.1 75.7 $\pm$ 2.0 87.1 $\pm$ 0.5 90.2 $\pm$ 0.3 92.1 $\pm$ 0.1 93.3 $\pm$ 0.2 94.0 $\pm$ 0.2 95.2 $\pm$ 0.1 95.6 $\pm$ 0.1
Herding [37] 20.2 $\pm$ 2.3 27.3 $\pm$ 1.5 34.8 $\pm$ 3.3 51.0 $\pm$ 3.1 63.5 $\pm$ 3.4 74.1 $\pm$ 2.5 80.1 $\pm$ 2.2 85.2 $\pm$ 0.9 88.0 $\pm$ 1.1 89.8 $\pm$ 0.9 94.6 $\pm$ 0.4 95.6 $\pm$ 0.1
k-Center Greedy [22] 18.5 $\pm$ 0.3 26.8 $\pm$ 1.2 31.1 $\pm$ 1.2 51.4 $\pm$ 2.1 75.8 $\pm$ 2.4 87.0 $\pm$ 0.3 90.9 $\pm$ 0.4 92.8 $\pm$ 0.1 93.9 $\pm$ 0.2 94.1 $\pm$ 0.1 95.4 $\pm$ 0.1 95.6 $\pm$ 0.1
Forgetting [47] 21.4 $\pm$ 0.5 29.8 $\pm$ 1.0 35.2 $\pm$ 1.6 52.1 $\pm$ 2.2 67.0 $\pm$ 1.5 86.6 $\pm$ 0.6 91.7 $\pm$ 0.3 93.5 $\pm$ 0.2 94.1 $\pm$ 0.1 94.6 $\pm$ 0.2 95.3 $\pm$ 0.1 95.6 $\pm$ 0.1
GraNd [50] 17.7 $\pm$ 1.0 24.0 $\pm$ 1.1 26.7 $\pm$ 1.3 39.8 $\pm$ 2.3 52.7 $\pm$ 1.9 78.2 $\pm$ 2.9 91.2 $\pm$ 0.7 93.7 $\pm$ 0.3 94.6 $\pm$ 0.1 95.0 $\pm$ 0.2 95.5 $\pm$ 0.2 95.6 $\pm$ 0.1
CAL [60] 22.7 $\pm$ 2.7 33.1 $\pm$ 2.3 37.8 $\pm$ 2.0 60.0 $\pm$ 1.4 71.8 $\pm$ 1.0 80.9 $\pm$ 1.1 86.0 $\pm$ 1.9 87.5 $\pm$ 0.8 89.4 $\pm$ 0.6 91.6 $\pm$ 0.9 94.7 $\pm$ 0.3 95.6 $\pm$ 0.1
DeepFool [27] 17.6 $\pm$ 0.4 22.4 $\pm$ 0.8 27.6 $\pm$ 2.2 42.6 $\pm$ 3.5 60.8 $\pm$ 2.5 83.0 $\pm$ 2.3 90.0 $\pm$ 0.7 93.1 $\pm$ 0.2 94.1 $\pm$ 0.1 94.8 $\pm$ 0.2 95.5 $\pm$ 0.1 95.6 $\pm$ 0.1
Craig [70] 22.5 $\pm$ 1.2 27.0 $\pm$ 0.7 31.7 $\pm$ 1.1 45.2 $\pm$ 2.9 60.2 $\pm$ 4.4 79.6 $\pm$ 3.1 88.4 $\pm$ 0.5 90.8 $\pm$ 1.4 93.3 $\pm$ 0.6 94.2 $\pm$ 0.2 95.5 $\pm$ 0.1 95.6 $\pm$ 0.1
GradMatch [71] 17.4 $\pm$ 1.7 25.6 $\pm$ 2.6 30.8 $\pm$ 1.0 47.2 $\pm$ 0.7 61.5 $\pm$ 2.4 79.9 $\pm$ 2.6 87.4 $\pm$ 2.0 90.4 $\pm$ 1.5 92.9 $\pm$ 0.6 93.2 $\pm$ 1.0 93.7 $\pm$ 0.5 95.6 $\pm$ 0.1
Glister [78] 19.5 $\pm$ 2.1 27.5 $\pm$ 1.4 32.9 $\pm$ 2.4 50.7 $\pm$ 1.5 66.3 $\pm$ 3.5 84.8 $\pm$ 0.9 90.9 $\pm$ 0.3 93.0 $\pm$ 0.2 94.0 $\pm$ 0.3 94.8 $\pm$ 0.2 95.6 $\pm$ 0.2 95.6 $\pm$ 0.1
Facility Location [97] 22.3 $\pm$ 2.0 31.6 $\pm$ 0.6 38.9 $\pm$ 1.4 60.8 $\pm$ 2.5 74.7 $\pm$ 1.3 85.6 $\pm$ 1.9 91.4 $\pm$ 0.4 93.2 $\pm$ 0.3 93.9 $\pm$ 0.2 94.5 $\pm$ 0.3 95.5 $\pm$ 0.2 95.6 $\pm$ 0.1
GraphCut [97] 24.3 $\pm$ 1.5 34.9 $\pm$ 2.3 42.8 $\pm$ 1.3 65.7 $\pm$ 1.2 76.6 $\pm$ 1.5 84.0 $\pm$ 0.5 87.8 $\pm$ 0.4 90.6 $\pm$ 0.3 93.2 $\pm$ 0.3 94.4 $\pm$ 0.3 95.4 $\pm$ 0.1 95.6 $\pm$ 0.1
TABLE III: Coreset selection robustness performances on CIFAR10 with 10% corrupted labels (i.e., 10% of total samples, randomly selected, have been reassigned to new class labels).
Fraction $(1-\alpha)$ 0.1% 0.5% 1% 5% 10% 20% 30% 40% 50% 60% 90% 100%
Random 20.4 $\pm$ 0.3 28.1 $\pm$ 1.8 34.6 $\pm$ 1.0 57.2 $\pm$ 1.5 67.8 $\pm$ 1.5 79.6 $\pm$ 1.2 84.5 $\pm$ 0.8 86.9 $\pm$ 0.3 88.2 $\pm$ 0.5 89.2 $\pm$ 0.2 90.6 $\pm$ 0.2 90.8 $\pm$ 0.1
-0.6 -2.7 -2.1 -7.3 -7.9 -7.5 -5.7 -5.2 -5.1 -4.8 -4.6 -4.8
Herding [37] 11.4 $\pm$ 0.9 10.8 $\pm$ 0.5 11.1 $\pm$ 0.9 10.6 $\pm$ 1.1 11.7 $\pm$ 0.9 26.0 $\pm$ 3.4 50.1 $\pm$ 1.3 71.0 $\pm$ 0.6 79.1 $\pm$ 1.8 84.6 $\pm$ 0.6 90.4 $\pm$ 0.1 90.8 $\pm$ 0.1
-8.8 -16.5 -23.7 -40.4 -51.8 -48.1 -30.0 -14.2 -8.9 -5.2 -4.2 -4.8
k-Center Greedy [22] 12.6 $\pm$ 1.3 14.3 $\pm$ 0.8 16.1 $\pm$ 1.0 29.6 $\pm$ 1.6 41.7 $\pm$ 3.0 62.0 $\pm$ 2.0 73.8 $\pm$ 1.8 80.2 $\pm$ 0.7 83.9 $\pm$ 0.7 86.6 $\pm$ 0.5 90.4 $\pm$ 0.3 90.8 $\pm$ 0.1
-5.9 -12.5 -15.0 -21.8 -34.1 -25.0 -17.1 -12.6 -10.0 -7.5 -5.0 -4.8
Forgetting [47] 21.8 $\pm$ 1.6 31.1 $\pm$ 1.0 35.0 $\pm$ 1.3 52.8 $\pm$ 1.2 66.4 $\pm$ 1.3 83.2 $\pm$ 1.0 88.9 $\pm$ 0.2 90.7 $\pm$ 0.2 91.0 $\pm$ 0.4 91.5 $\pm$ 0.4 90.9 $\pm$ 0.1 90.8 $\pm$ 0.1
+0.4 +1.3 -0.2 +0.7 -0.6 -3.4 -2.8 -2.8 -3.1 -3.1 -4.4 -4.8
GraNd [50] 11.5 $\pm$ 0.9 11.9 $\pm$ 0.8 11.1 $\pm$ 0.6 10.8 $\pm$ 1.1 10.6 $\pm$ 1.2 25.4 $\pm$ 0.9 44.8 $\pm$ 2.0 67.2 $\pm$ 2.6 79.4 $\pm$ 1.3 86.3 $\pm$ 0.3 90.2 $\pm$ 0.2 90.8 $\pm$ 0.1
-6.2 -12.1 -15.6 -29.0 -42.1 -52.8 -46.4 -26.5 -15.2 -8.7 -5.3 -4.8
CAL [60] 21.3 $\pm$ 1.7 30.8 $\pm$ 1.0 36.8 $\pm$ 1.3 59.9 $\pm$ 0.8 71.3 $\pm$ 1.0 80.0 $\pm$ 0.2 83.9 $\pm$ 0.6 87.1 $\pm$ 0.3 89.1 $\pm$ 0.2 90.6 $\pm$ 0.2 91.9 $\pm$ 0.1 90.8 $\pm$ 0.1
-1.4 -2.3 -1.0 -0.1 -0.5 -0.9 -2.1 -0.4 -0.3 -1.0 -2.8 -4.8
DeepFool [27] 17.4 $\pm$ 0.9 21.6 $\pm$ 1.4 25.3 $\pm$ 1.3 33.6 $\pm$ 0.4 43.9 $\pm$ 3.2 65.5 $\pm$ 1.6 77.0 $\pm$ 1.3 84.5 $\pm$ 0.5 86.6 $\pm$ 0.9 88.7 $\pm$ 0.5 90.8 $\pm$ 0.2 90.8 $\pm$ 0.1
-0.2 -0.8 -2.3 -9.0 -16.9 -17.5 -13.0 -8.6 -7.5 -6.1 -4.7 -4.8
Craig [70] 19.5 $\pm$ 1.4 20.2 $\pm$ 1.4 24.8 $\pm$ 1.1 30.4 $\pm$ 0.9 31.7 $\pm$ 1.6 39.2 $\pm$ 1.5 58.4 $\pm$ 2.9 73.1 $\pm$ 1.4 81.4 $\pm$ 0.6 85.3 $\pm$ 0.4 90.5 $\pm$ 0.3 90.8 $\pm$ 0.1
-3.0 -6.8 -6.9 -14.8 -28.5 -40.4 -30.0 -17.7 -11.9 -8.9 -5.0 -4.8
GradMatch [71] 15.7 $\pm$ 2.0 21.4 $\pm$ 0.7 23.0 $\pm$ 1.6 27.6 $\pm$ 2.4 31.4 $\pm$ 2.5 37.7 $\pm$ 2.1 55.6 $\pm$ 2.5 72.0 $\pm$ 1.4 80.3 $\pm$ 0.4 85.3 $\pm$ 0.5 90.2 $\pm$ 0.2 90.8 $\pm$ 0.1
-1.7 -4.2 -7.8 -19.6 -30.1 -42.2 -31.8 -18.4 -12.6 -7.9 -3.5 -4.8
Glister [78] 14.9 $\pm$ 2.0 20.9 $\pm$ 1.5 24.5 $\pm$ 1.3 29.0 $\pm$ 1.9 31.7 $\pm$ 2.1 40.2 $\pm$ 2.3 57.2 $\pm$ 1.3 72.2 $\pm$ 1.3 80.6 $\pm$ 0.4 85.4 $\pm$ 0.5 90.2 $\pm$ 0.2 90.8 $\pm$ 0.1
-4.6 -6.6 -8.4 -21.7 -34.6 -44.6 -33.7 -20.8 -13.4 -9.4 -5.4 -4.8
Facility Location [97] 20.7 $\pm$ 1.4 24.5 $\pm$ 0.8 29.5 $\pm$ 1.2 43.1 $\pm$ 2.9 51.3 $\pm$ 1.7 61.5 $\pm$ 2.0 75.8 $\pm$ 0.6 81.5 $\pm$ 0.4 84.9 $\pm$ 0.3 87.2 $\pm$ 0.3 90.6 $\pm$ 0.2 90.8 $\pm$ 0.1
-1.6 -7.1 -9.4 -17.7 -23.4 -24.1 -15.6 -11.7 -9.0 -7.3 -4.9 -4.8
GraphCut [97] 22.0 $\pm$ 0.7 34.0 $\pm$ 2.2 41.3 $\pm$ 1.2 66.0 $\pm$ 1.2 77.5 $\pm$ 0.6 84.6 $\pm$ 0.5 87.6 $\pm$ 0.2 89.8 $\pm$ 0.4 87.8 $\pm$ 0.2 87.2 $\pm$ 0.1 90.0 $\pm$ 0.2 90.8 $\pm$ 0.1
-2.3 -0.9 -1.5 +0.3 +0.9 +0.6 -0.2 -0.8 -5.4 -7.2 -5.4 -4.8
VII Applications
Coreset selection has found applications in various domains of machine learning, spanning computer vision [98, 74, 99, 100, 101], natural language processing [102, 103, 104, 105], and quantum machine learning [106, 107, 108]. In the following sections, we examine specific fields where coreset selection has made interesting contributions.
VII-A Image Enhancement
Coreset selection has also proven valuable in image enhancement tasks, specifically in the domain of image super-resolution, where practitioners often construct training sets by randomly cropping patches from high-resolution images (800 2K images lead to 32K training images) [109, 110, 101]. Although simple, this patch-based extraction frequently yields a large proportion of redundant low-frequency patches that convey limited additional information [111, 112].
By adopting a score-based training-oriented coreset selection strategy, one that ranks patches based on their importance to the model’s learning dynamics (i.e., Sobel filter or loss value-based), it becomes possible to discard uninformative patches in favor of fewer but more critical (high-frequency) examples. This not only leads to a substantial reduction in dataset size and training overhead, but recent empirical results further show that such curated training sets can improve the reconstruction quality of super-resolution networks.
In addition, we can use score-based patch ranking to weight image patches for deeper refinement during training, as proposed by ClassSR [113] and others [114, 115, 116, 117].
VII-B Neural Architecture Search
Neural Architecture Search (NAS) automates the design of network topologies and sometimes outperforms manually engineered architectures at the cost of substantial computational overhead [118, 119]. A key bottleneck arises when evaluating candidate architectures on large datasets, as each candidate in the search space must be trained (or partially trained) to assess its fitness [120, 121]. This computational burden grows significantly when the dataset itself is extensive [122, 123].
Recent work shows that coreset selection can overcome these scalability issues by constructing coresets, i.e., smaller yet sufficiently representative subsets used during the architecture search phase [124, 125, 126, 127]. For instance, ranking and filtering out less informative training samples often shortens the NAS procedure by a factor of two or more while simultaneously improving the resulting architecture’s accuracy once retrained on the full dataset.
VII-C Dataset Distillation
Coreset selection and dataset distillation share the same high-level goal: reducing the size of a dataset. Yet, they differ in the nature of the compressed data. In dataset distillation, one synthesizes entirely new samples that preserve training quality [128, 129, 130].
Concretely, let the original image classification dataset be $\mathcal{T}=(X_{r},Y_{r})$ with $N$ real examples. Dataset distillation tries to generate a small, synthetic set $\mathcal{S}=(X_{s},Y_{s})$ of $M\ll N$ synthetic samples. We define this objective as
$\mathcal{S}^{*}=\arg\min_{\mathcal{S}}\mathcal{L}\bigl{(}\mathcal{S},\mathcal{T}\bigr{)},$ (46)
where $\mathcal{L}$ measures how good the synthetic set $\mathcal{S}$ represents the original dataset $\mathcal{T}$. Concrete implementations of $\mathcal{L}$ are, for example, methods like Dataset Condensation [131], Distribution Matching [132], or Matching Training Trajectories [133] and others [134, 135].
Since both coreset selection and dataset distillation ultimately aim for the same destination, it is only natural to explore how coreset and distillation methods might be combined for even more powerful dataset reduction [136, 137, 138].
For instance, He et al. [139] apply coreset selection directly to the distilled dataset by combining a logit-based prediction-error criterion—aiming to retain “easy” samples—with a Rademacher-complexity measure that maintains class balance. Separately, Xu et al. [140] propose a complementary pipeline in which coreset selection, guided by empirical loss, prunes real data before the distillation stage and is further enhanced by a dynamic pruning mechanism during the distillation process.
VII-D Imbalanced Datasets
Class imbalance poses significant challenges for machine learning models, often leading to biased predictions toward majority classes [141, 142]. Coreset selection techniques have been explored as a potential solution to this problem [143].
For instance, in federated learning, where multiple clients collaborate in training together, class imbalance is a major issue. It has been shown that adapting the principles of coreset selection to this scenario can mitigate the imbalance problem [144, 145]. As a result, coordinating coreset selection across clients achieves a globally balanced dataset while respecting the privacy constraints necessary in federated settings.
VII-E Continual Learning
Continual Learning (CL) aims to enable machine learning models to learn new tasks while preserving previously acquired knowledge incrementally [146, 147]. A fundamental challenge in CL is catastrophic forgetting, where a model forgets prior knowledge when trained on new data [148]. For instance, a model could encounter classes in the current tasks that are not appearing in later tasks, which results in forgetting them. Rehearsal-based strategies, which maintain a subset of past examples for replay during training, have emerged as an effective solution to mitigate forgetting [149, 150].
However, storing and retraining on large amounts of past data is computationally expensive and memory-intensive. Coreset selection techniques play a crucial role in optimizing the selection of representative samples for rehearsal, ensuring efficient learning while reducing memory overhead [151, 152, 153].
VII-F Reinforcement Learning
Similar to CL, Reinforcement Learning (RL) requires extensive memory sampling to not forget past experiences [154]. As RL is naturally operating in interactive environments (e.g., in robotics), retraining agents also in parallel on past experiences is costly in terms of both data and computation [155]. Thus, coreset selection has been used to improve experience replay by selecting the most informative past experiences for training [156, 157].
VII-G Machine Unlearning
Machine unlearning refers to the process of removing the influence of specific data points from a trained machine learning model without retraining it. This is particularly useful in scenarios where user privacy concerns require the deletion of data, or when models must “forget” outdated or incorrect information.
Coreset selection can play a crucial role in this process by identifying the most important data points to either retain or remove, thereby minimizing the impact on the model’s overall performance while improving the cost of unlearning [158]. Utility Preserving Coreset Selection (UPCORE) is a framework introduced by [159] for data selection that minimizes the performance degradation during unlearning, focusing on which data points to remove (i.e, the ”forget set”) but also leveraging a coreset selection algorithm to identify which samples to retain.
VII-H Self-supervised Learning
Using coresets in self-supervised learning can improve the efficiency and effectiveness of SSL pre-training. This is critical when training on large open sets (e.g., ImageNet) for domain-specific tasks (e.g., fine-grained classification) where distribution mismatch between the open set and target dataset can degrade performance. Kim et al. [160] propose Simcore, a technique that samples a coreset that minimizes the latent-space distance between the open-set and target dataset. Singh et al. [161] utilize Bloom filters for rapid retrieval of the coreset from the open set and integrate it with Simcore to achieve a reduction in sampling time. RETRIEVE [76] achieves a speedup of around 3$\times$ in the traditional SSL setting.
VII-I Large Language Models
Using coresets in self-supervised learning (SSL) enhances both the efficiency and effectiveness of pre-training, particularly when dealing with large open sets (e.g., ImageNet) for domain-specific tasks such as fine-grained classification. In such scenarios, a distribution mismatch between the open set and the target dataset can negatively impact performance. Kim et al. [160] introduce Simcore, a method that selects a coreset to minimize the latent-space distance between the open-set and target dataset. Similarly, RETRIEVE [76] accelerates traditional SSL training by approximately 3× speedup.
VIII Discussion and Future Directions
Coreset selection has evolved significantly beyond its early geometric origins, fueled by deep learning’s growing scale and resource needs. As a result, coreset selection finds itself applied to many machine learning tasks. As surveyed, training-free methods remain conceptually straightforward and computationally light, yet often underperform in comparison to training-oriented methods in tasks demanding fine-grained or model-specific subsets.
On the other hand, training-based approaches, particularly those relying on loss signals, gradient information, and submodular formulations, can yield competitive or even state-of-the-art performance, albeit at a higher computational cost. The recent development in active learning and blind coreset selection further broadens the field, approaching fully unlabeled environments.
VIII-A Discussion
Despite these advancements, we observe the following:
Trade-offs Between Simplicity and Specificity. Training-free methods (e.g., random, Herding, and k-Center Greedy) perform surprisingly well in some scenarios and scale effortlessly to massive datasets. By contrast, training-oriented or bilevel approaches (e.g., GradMatch, GLISTER) finely tune subsets for a specific architecture or objective but can be sensitive to random seeds, hyperparameters, and model changes.
Coverage vs. Difficulty. Methods that prioritize “hard” samples often improve decision boundaries but risk ignoring “easy” or “redundant” data that might be crucial for overall distribution coverage. Future work may consider hybrid approaches that adaptively select from both easy and hard examples, especially in dynamic or incremental learning settings, similar to Zheng et al. [10].
Blind or Label-Free Approaches. Label-free frameworks like ELFS and ZCore offer a promising alternative for scenarios where supervision is limited or nonexistent. However, the quality of their derived coresets hinges on the quality of pre-trained encoders (or pseudo-labeling), which may degrade under domain shifts or for highly specialized tasks.
Influence of Dataset Size and Pruning Ratio. As indicated by recent theoretical and empirical studies, the optimal coreset selection strategy may depend on whether $\mathcal{T}$ is small ($\sim\!1\mathrm{k}$$10\mathrm{k}$ samples) or extremely large ($\sim\!1\mathrm{M}$$10\mathrm{M}$). While large datasets often contain heavy redundancy, smaller ones might demand strategies that retain more “easy” or high-density examples for stable generalization. The same paradigm shift goes for low and high pruning ratios.
Evaluation and Benchmarking. Different works evaluate coreset selection on varying metrics and apply different post-processing (e.g., fine-tuning for 100 vs. 300 epochs). This complicates direct comparisons. Comprehensive benchmarks, possibly spanning multiple architectures and data modalities (image, text, speech), remain an open challenge, while DeepCore [15] offers an initial step in this direction.
Summary. For high pruning ratios, submodular function maximization seems a good strategy, effectively selecting a diverse and representative subset of the data. However, for low pruning ratios, score-based methods are preferable due to their efficiency in retaining the most informative samples without excessive computational overhead (e.g., forgetting). That said, submodular approaches must be used with caution, as they can become computationally intractable if the resulting coreset remains large, even at high pruning ratios. Balancing selection quality with computational feasibility remains a key challenge in practical applications.
VIII-B Future Research Directions
In this section, we want to give a short summary of possible research avenues that can be explored in the future.
Adaptive and Continual Coresets. An important direction is continual learning or streaming scenarios, where new data arrives incrementally. Approaches must update the coreset without full retraining or batch-level re-computation. Combining online gradient matching or submodular updates with incremental embedding updates is a rich research frontier.
Fairness and Robustnes. Beyond mere accuracy, future coreset research can incorporate fairness constraints (ensuring demographic representation) or robust objectives (countering adversarial or noisy labels). Submodular or bilevel methods could be extended to optimize these multicriteria.
Self-Evolving Coresets with Meta-Learning. A highly ambitious direction is to design meta-learning frameworks that continuously refine their subset selection strategy based on the data’s evolving characteristics. Instead of defining a fixed scoring or submodular criterion from the outset, the system would learn how to pick coresets for new tasks or distributions by observing the performance and training dynamics on prior tasks.
Multi-Modal Data. Most coreset research focuses on single-modality data (e.g., images). Extending coverage or gradient-based selection to multi-modal tasks (e.g., image and text) is non-trivial, as each modality’s representation must be jointly considered.
IX Conclusion
This survey has traced the evolution of coreset selection from its geometric foundations to more sophisticated training-oriented and blind methods. By examining core approaches - random and geometry-based sampling, gradient matching, submodular formulations, and bilevel optimization - this work reveals how different strategies strike varied balances between computational efficiency, label requirements, and approximation accuracy.
From a birds-eye view, we highlight how both small and large datasets can be pruned effectively, though the preferred criteria may shift as sample sizes grow or task complexities change. Furthermore, emerging perspectives on coverage, margin-based difficulty, and meta-learning provide broader frameworks for understanding what genuinely constitutes the most “informative” data.
Taken together, these developments underscore how data pruning is not simply about efficiency but can boost generalization by preserving the crucial patterns of the original dataset. As coreset selection continues to mature, it stands as a promising research direction to mitigate the resource demands of modern deep learning, adapt to new domains where little to no supervision is available, and potentially unify competing priorities such as fairness, robustness, and interpretability.